A particle has a charge of q 3e, where e is the charge on an electron. (a) Determine the electric potential (in V) due to the charge at a distance r 0.260 cm from the charge. (b) Determine the electric potential difference (in V) between a point that is 3r away and this point, that is V(3r) - r) (c) Determine the electric potential difference (in V) between a point that is-away and this point, that is V(-)-V(r) (d) How would the answers change if the electrons are replaced by protons? O The sign of answer (a) would change O The sign of answer (b) would change. O The sign of answer (e) would change O The sign of all answers would change. All answers would remain the same.
Answer #1

V = - kq/r = 9× 10^9 × 3×1.6 ×10^-19/ 0.26×10^-2 = 166.15 X 10^-8 = - 1. 66×10^-6 V

V ( 3r) = - 0.554 x10^-6 + 1.66 × 10^-6 = 1.11 x10^-6 V

V (r/3) = - 4.98 X 10^-6 + 1.66 ×10^-6 = - 3.32×10^-6V

The sign of all answers would change